Gate2018 cs Q47

  • Option : A
  • Explanation : The condition probability table given is
     HDMDLD
    HG0.400.480.12
    MG0.100.650.25
    LG0.010.500.49

    P(HG) = 0.2
    P(MG) = 0.5
    P(LG) = 0.3
    Drawing the tree diagram for HD we get,

    From diagram, P(HG ∩ HD) = 0.2x0.4
    P(HD) = 0.2x0.4+0.5x0.1+0.3x0.01 = 0.133
    Required probability,
    P(HG | HD) = (0.2x0.4)/0.133 = 0.60
Cancel reply

Your email address will not be published. Required fields are marked *


Cancel reply

Your email address will not be published. Required fields are marked *